The simplified consolidate balance sheet shown below is for the entire commercial banking system. All figures are in billions. The reserve ratio is 20 percent. Assets Liabilities and Net Worth Vault Cash 25,000 Checkable Deposits 220,000 Reserves Deposits 65,000 Shareholder’s Equity 80,000 Loans 115,000 Securities 45,000 Property 50,000 a. Calculate the actual reserves: ______________ b. Calculate the required reserves: ______________ c. Calculate the excess reserves: ______________ d. What is the money multiplier? _____________ e. What is the maximum amount of additional loans the banking system can create?
Answer #1

a) Actual reserves = 65000

b) Required reserves = 20% of Checkable deposits = 20% of 220,000 = 44,000

c) Excess reserves = Actual reserve - Required reserve = 65000 - 44000 = 21000

d) Multiplier = 1/Required reserve = 1/0.20 = 5

e) Banks can give loan of excess reserves i.e. $ 21000.